# Fight like a Woman:

### v1

@@ Introduction
If the gender conviction rate gap is stronger in cases where the defendant was not caught red-handed, then this may be because female and male judges interpret more subjective pieces of evidence differently. This would provide possible explanatory force for the observed leniency gap in domestic violence cases. If the gender conviction rate gap is similar across domestic violence cases regardless of the type of evidence presented, then there must be other forces driving this lenience gap.
@@
Quarter-District FE YES YES YES YES YES YES
Judge Career Controls YES YES YES YES YES YES
@@
We find that a gender conviction rate gap exists only in those cases where the defendant was not caught red-handed. We only find significant results for the female coefficient ($\beta$) in Column (2) and the interaction coefficient ($\beta_{3}$) in Column (6). These results suggest that female and male judges interpret less objective evidence differently, while they interpret objective evidence (cases flagged as in flagrante delicto) in similar ways. These different interpretations of more subjective cases offer one explanation for the gender conviction rate gap in domestic violence cases.
@@
Next, we analyze the association between the strength of career incentives and the size of the conviction gender gap. We show that male and female judges differ more intensely at the end of their careers. This corresponds with the period where they face the lowest probability of upcoming performance evaluation for promotion to the Appeals Court. [^13] See Appendix B for more information on the proportion of cases analyzed by judges who apply for promotion in the near future Career pressures may lead early-career male and female judges to act more similarly, possibly in response to feedback from Appeals Court magistrates, who are overwhelmingly male.
@@
Quarter-District FE YES YES YES YES YES YES
Case Controls YES YES YES YES YES YES
@@ Conclusion
In this paper, we provide evidence that cases assigned to female judges are significantly more likely to result in conviction than those assigned to a male judges with a similar career path. This pattern is, to a large extent, due to the reasons that are specific to domestic violence. The gender difference in conviction rates is stronger for cases involving domestic violence than for crimes that are similar in nature (physical assault) and sentence length (misdemeanors).
Furthermore, we find suggestive evidence for two major drivers of the conviction gender gap in domestic violence. The first is that female judges are stricter than men in cases of intimate partner violence but equally strict for cases of other domestic violence offenses. We claim this is evidence of identification with one’s own gender, as perception of intimate partner violence is highly affected by patriarchal social norms, as opposed to other forms of violence where this might not be the case.
The second driver comes from gender differences in the way judges analyze evidence. There is no gender gap when the defendant is caught red-handed by the police, making the evidence presented clear and objective, but the gap is very strong when there is more room for subjectivity in the analysis.
@@ Appendix B Final-stage Judges Face Weaker Career Incentives
Quarter-District FE YES YES YES YES
@@
Quarter-District FE YES YES YES YES

### v2

@@ Introduction
If the gender conviction rate gap is stronger in cases where the defendant was not caught red-handed, then this may be because female and male judges interpret more subjective pieces of evidence differently (informational account). If the gender conviction rate gap is similar across domestic violence cases regardless of the type of evidence presented, then there must be other forces driving this lenience gap.
@@
Quarter-District FE $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Judge Career Controls $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@
We find that a gender conviction rate gap exists only in those cases where the defendant was not caught red-handed. We only find significant results for the female coefficient ($\beta$) in Column (2) and the interaction coefficient ($\beta_{3}$) in Column (6). These results suggest that female and male judges interpret less objective evidence differently, while they interpret objective evidence (cases flagged as in flagrante delicto) in similar ways. These different interpretations of more subjective cases are consistent with the informational account being one possible driving force of the in-group bias documented in Section 4.
@@
In this section, we analyze the association between the strength of career incentives and the size of the conviction gender gap. We show that male and female judges differ more intensely at the end of their careers. This corresponds with the period where they face the lowest probability of upcoming performance evaluation for promotion to the Appeals Court. [^13] See Appendix B for more information on the proportion of cases analyzed by judges who apply for promotion in the near future Career pressures may lead early-career male and female judges to act more similarly, possibly in response to feedback from Appeals Court magistrates, who are overwhelmingly male (Laneuville, 2024).
@@
Quarter-District FE $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Case Controls $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@ Conclusion
In this paper, we provide evidence that cases assigned to female judges are significantly more likely to result in conviction than those assigned to male judges with similar career paths. This pattern is, to a large extent, due to the reasons that are specific to domestic violence. The gender difference in conviction rates is stronger for cases involving domestic violence than for crimes that are similar in nature (physical assault) and sentence length (misdemeanors), suggesting the existence of a type of in-group bias as defined by Shayo and Zussman (2011).
Furthermore, we find suggestive evidence for two major drivers of the conviction gender gap in domestic violence. The first is that female judges are stricter than men in cases of intimate partner violence but equally strict for cases of other domestic violence offenses. We claim this is evidence of identification with one’s own gender, as perception of intimate partner violence is highly affected by patriarchal social norms, as opposed to other forms of violence where this might not be the case. This result is consistent with the representational account as defined by Boyd, Epstein, and Martin (2010).
The second driver comes from gender differences in the way judges analyze evidence. There is no gender gap when the defendant is caught red-handed by the police, making the evidence presented clear and objective, but the gap is very strong when there is more room for subjectivity in the analysis. This result is consistent with the informational account as defined by Boyd, Epstein, and Martin (2010).
@@ Appendix B Final-stage Judges Face Weaker Career Incentives
Quarter-District FE $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@
Quarter-District FE $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@
Appendix C Modelling the Gender Conviction-Rate Gap
In this appendix, we propose a simple economic model to formalize the concepts of the gender conviction-rate gap, in-group bias (Shayo and Zussman, 2011) and their driving forces, the representational and informational accounts (Boyd, Epstein, and Martin, 2010). First, to formally define the first two concepts, we adapt the potential outcome model (Rubin, 1974) to a scenario where cases of different types of crime may be treated with a female judge instead of a male judge. Second, to formally define the representational and informational accounts of the in-group bias, we adapt the threshold-crossing model (Heckman and Vytlacil, 2005) to a context where the latent heterogeneity is associated with the evidence of a criminal case and the propensity score is associated with the judge’s conviction criteria.
We start by looking at criminal cases as our observational unit. Each case is associated with a type of crime: $d=1$ indicates that it is a domestic violence offense, and $d=0$ indicates that it is either a misdemeanor offense or another type of physical assault case. Moreover, each case may receive two randomly allocated treatments: $f=1$ indicates that the case was assigned to a female judge and $f=0$ indicates that the case was assigned to a male judge. Finally, each case has a potential conviction decision, $Y\left(f,d\right)$, which depends on the gender of the presiding judge and on the type of crime. [^16] Note that, in our context, sex is a manipulable variable because the unit of observation is the case instead of the judge. Moreover, type of crime is a pre-determined covariate, because it is chosen by the prosecutor before the presiding judge is assigned.
We define the gender conviction-rate gap as
$\mathbb{E}\left[Y\left(1,1\right)-Y\left(0,1\right)\right]\neq 0,$ (C.1)
i.e., the probability that a domestic violence case ends in a conviction when presided by a female judge is different from the probability that a domestic violence case ends in a conviction when presided by a male judge.
We also define the in-group bias as
$\left\{\begin{matrix}[l]\mathbb{E}\left[Y\left(1,1\right)-Y\left(0,1\right)\right]&>&0\\ \mathbb{E}\left[Y\left(1,1\right)-Y\left(0,1\right)\right]&>&\mathbb{E}\left[Y\left(1,0\right)-Y\left(0,0\right)\right]\end{matrix}\right.,$ (C.2)
i.e., the gender conviction-rate gap for domestic violence cases (i) is positive, and (ii) is larger than the gender conviction-rate gap for other types of crime.
Next, we formally define the representational and informational account using our empirical context and a threshold-crossing model. In any criminal case, evidence is produced against the defendant. In our model, the judge sets a conviction criterion and interprets the evidence. Then, the judge convicts the defendant if the judge interprets the evidence as being sufficient to do so.
The evidence interpretation function, $E\colon\left\{0,1\right\}^{3}\times\mathbb{R}\rightarrow\mathbb{R}$, depends on the judge’s gender (a female indicator variable, $F$), the type of crime (a domestic violence indicator variable, $D$), whether there is objective and strong evidence against the defendant (an indicator variable $R$ that is equal to one if the defendant was caught red-handed when committing the crime), and all other types of evidence produced by the police and the prosecution, $V$. Hence, $E\left(F,D,R,V\right)$ is a random variable capturing the amount of evidence against the defendant as interpreted by the judge assigned to the case.
The judge’s conviction criterion function, $P\colon\left\{0,1\right\}^{3}\rightarrow\mathbb{R}$, depends on the judge’s gender, the type of crime and whether the crime is associated with intimate partner violence (an indicator variable $W$ that is equal to one if the victim is the wife or intimate partner of the defendant). Hence, $P\left(F,D,W\right)$ is a random variable capturing the conviction criterion of the judged assigned to the case.
The judge convicts the defendant if the judge interprets the evidence as being sufficient to do so, i.e.,
$Y=\mathbf{1}\left\{P\left(F,D,W\right)\leq E\left(F,D,R,V\right)\right\},$ (C.3)
where $Y$ denotes whether the defendant is convicted or not.
Note that Equation (C.3) imposes two exclusion restrictions. While the relationship status between the victim and the defendant — $W$ —- only enters the judge’s conviction criterion function, being caught red-handed — $R$ — only enters the evidence interpretation function. The first exclusion restriction is imposed because the Criminal Process Code does not include this variable as valid evidence against the defendant, but the Criminal Law Code includes this variable as an aggravating factor when setting the sentence. The second exclusion restriction is imposed because the Criminal Process Code considers in flagrante delicto as a piece of evidence against the defendant but not as an aggravating factor.
Using this model, we can formally define the representational account driving the in-group bias as
$P\left(1,1,1\right)-P\left(0,1,1\right)<P\left(1,1,0\right)-P\left(0,1,0\right)\leq 0,$ (C.4)
i.e., female judges may set stricter conviction criteria than male judges for domestic violence cases and female judges are even stricter when the case is associated with intimate partner violence.
We can also formally define the informational account driving the in-group bias as
$E\left(1,1,0,v\right)-E\left(0,1,0,v\right)>E\left(1,1,1,v\right)-E\left(0,1,1,v\right)\geq 0,$ (C.5)
for any $v$ in the support of $V$. When these inequalities hold, the female judge interprets the evidence more harshly than a male judge when the defendant is caught red-handed and the female judge interprets the evidence even more harshly when the defendant is not caught red-handed.
In the main text, we use data from criminal offenses in the State of Sao Paulo, Brazil, to test the existence of each one of the four concepts defined above. In Table 3, we show that the gender conviction-rate gap (Equation (C.1)) exists. Then, in Table 4, we show that in-group bias (Equation (C.2)) is present in our application. Finally, Table 5 shows that our data is consistent with the representational account (Equation (C.4) while Table 6 provides evidence that our data is consistent with the informational account (Equation (C.5).